OpenStudy (anonymous):

find the distance between the points G(3,6) and H (4,7) ? Distance GH= ? 1 dec.pl.

OpenStudy (anonymous):

1.4

jimthompson5910 (jim_thompson5910):

d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((4-3)^2+(7-6)^2) d = sqrt((1)^2+(1)^2) d = sqrt(1+1) d = sqrt(2) d = 1.414 So distance between the two points is approximately 1.4 units.